# On the universal curve with unordered marked points in positive characteristic

### v2

@@ 1. Introduction
Here $\pi_{1}(C,\bar{y})^{{(\ell)}}$ denotes the pro-$\ell$ completion of $\pi_{1}(C,\bar{y})$, and $\pi_{1}^{\prime}\!\bigl({\mathcal{C}}_{g,[n]/\bar{k}},\bar{y}\bigr)$ is a canonical quotient of $\pi_{1}\!\bigl({\mathcal{C}}_{g,[n]/\bar{k}},\bar{y}\bigr)$ obtained by the SGA 1 construction (cf. [SGA1, Exp. XIII, §4]). We briefly review this construction in §2.1.
@@
This paper is a sequel to [LW25], where the corresponding nonsplitting results for universal curves with unordered marked points were established in characteristic 0. Our aim here is to extend those results to positive characteristic. Although the overall strategy follows the same general framework as in the characteristic 0 case, carrying out this extension requires substantial additional work. In characteristic 0, the essential geometric information needed for the argument is encoded in the weight filtration on relative and continuous relative completions. In positive characteristic, while continuous relative completion remains available for algebraic fundamental groups, transferring the corresponding structural properties across characteristics is far from automatic. To address this, we make systematic use of relative pro-$\ell$ completion and continuous relative completion, together with comparison isomorphisms between characteristic 0 and characteristic $p$ established in [wat_rational_points_inp], which allow us to transport the necessary monodromy and homotopy-theoretic information into characteristic $p$. These tools play a central role throughout the paper and are indispensable for the arguments that follow. Familiarity with the methods developed in [LW25] will therefore be beneficial to the reader.
@@ 2. Moduli of Curves with unordered marked points
The symmetric group $S_{n}$, regarded as a finite constant group scheme over $\operatorname{Spec}{\mathbb{Z}}$, is finite étale. By base change,
@@
is a finite constant étale group scheme over $k$ acting on ${\mathcal{M}}_{g,n/k}$ by permuting the marked points. Since ${\mathcal{M}}_{g,n/k}$ is a smooth Deligne–Mumford stack and $S_{n/k}$ is a constant finite étale group scheme acting on it, the quotient stack
@@
is again a smooth Deligne–Mumford stack over $k$. It parametrizes smooth, proper genus $g$ curves with $n$ unordered marked points.
@@ 2.1. Homotopy exact sequences in positive characteristic
In this subsection, we briefly recall the SGA 1 construction of the homotopy exact sequence for a proper smooth morphism with geometrically connected fibers, and then apply it to the ordered and unordered universal curves (cf. [SGA1, SGA 1, Exp. XIII, §4]; see also [wat_rational_points_inp, Prop. 4.2]).
Let $L$ be a field, and let $T$ be a connected, locally noetherian Deligne–Mumford stack over $L$. Let $f:C\longrightarrow T$ be a proper smooth morphism whose geometric fibers are connected curves of genus $g\geq 2$. Fix geometric points $\bar{\zeta}$ in $T$ and $\bar{a}$ in the fiber $C_{\bar{\zeta}}$ of $f$ over $\bar{\zeta}$. Let $\ell$ be a prime number different from $\operatorname{char}(k)$. Following [SGA1, SGA 1, Exp. XIII, §4.3–4.4], set
@@
We also write $\pi_{1}(C_{\bar{\zeta}},\bar{a})^{(\ell)}$ for the maximal pro-$\ell$ quotient of $\pi_{1}(C_{\bar{\zeta}},\bar{a})$. With this notation, [wat_rational_points_inp, Prop. 4.2] yields the exact sequence
@@ 3.1. The pro-$\ell$ completion of $\pi_{1}(F(C,n))$
Let $K$ be the fraction field of $R$ and $\bar{K}$ an algebraic closure of $K$. For an integer $n\geq 1$, let $F(X,n)$ denote the ordered configuration space of $X$ with $n$ labeled points. The Fulton–MacPherson compactification [FM] of $F(X,n)$ is proper and smooth over $R$, and $F(X,n)$ is the complement of a relative normal crossings divisor.
@@
for each prime $\ell\neq p$. By [SGA1, Exposé XIII, Thm. 2.10, Cor. 2.9] and the fact that the Fulton–MacPherson compactification is smooth over $R$ and that $F(X,n)$ is the complement of a relative normal crossings divisor, this specialization map $\mathrm{sp}_{\ell}$ is an isomorphism.
@@ 3.2. The $\ell$-adic unipotent completion
We now pass to the $\ell$-adic unipotent completions. Following [hain-matsumoto, Appendix A], let $F_{g,n}^{\mathrm{un}}$ and $F_{g,[n]}^{\mathrm{un}}$ denote the $\ell$-adic unipotent completions of $F_{g,n}$ and $F_{g,[n]}$ over $\mathbb{Q}$$_{\ell}$, respectively. These are pro-unipotent algebraic $\mathbb{Q}$$_{\ell}$}-groups equipped with Zariski-dense continuous homomorphisms
@@ 4. Relative Pro-$\ell$ Completion
To overcome this, one instead considers the relative pro-$\ell$ completion. This construction retains essential information about the mapping class group by taking into account the natural $\ell$-adic monodromy representation $\rho:{\Gamma}_{g}\longrightarrow\operatorname{Sp}_{2g}({\mathbb{Z}}_{\ell})$ and completing ${\Gamma}_{g}$ in a manner that is compatible with $\rho$. In this way, the resulting relative pro-$\ell$ completion preserves the $\ell$-adic structure coming from the geometry of ${\mathcal{M}}_{g}$. As shown in [HainMatsumoto2009], the resulting relative pro-$\ell$ completion is nontrivial and captures the geometric structure of ${\mathcal{M}}_{g}$ much more effectively than the ordinary pro-$\ell$ completion. For instance, when marked points are present, the mapping class group embeds into its relative pro-$\ell$ completion ([HainMatsumoto2009, Prop. 3.1]).
Because the second author applied this approach in the positive characteristic setting in [wat_rational_points_inp], we briefly recall below the definition and the basic properties of relative pro-$\ell$ completion. Full details and extensive background can be found in [HainMatsumoto2009].
@@ 4.1. The relative pro-$\ell$ completion of the mapping class group
When $\operatorname{char}(L)=p>0$ and $\ell\neq p$, the second author proved in [wat_rational_points_inp, Thm. 5.6] that the relative pro-$\ell$ completion of $\pi_{1}({\mathcal{M}}_{g,n/\bar{L}},\bar{\zeta})$ with respect to $\rho_{{\mathbb{Z}}_{\ell},\bar{\zeta}}$ is naturally isomorphic to ${\Gamma}_{g,n}^{{\mathrm{rel}(\ell)}}$, the relative pro-$\ell$ completion of the mapping class group ${\Gamma}_{g,n}$ with respect to its standard symplectic representation
@@ 4.2. Specialization for the universal curve in the ordered case
The rows are exact by right exactness of relative pro-$\ell$ completion and the center-freeness of $\pi_{1}(C_{\bar{\xi}})^{{(\ell)}}(\cong\pi_{1}(C_{\bar{\eta}})^{{(\ell)}})$, and the right vertical homomorphisms are isomorphisms (see the proof of [wat_rational_points_inp, Thm. 5.6]). It therefore follows that all middle vertical homomorphisms are isomorphisms. ∎
@@ 4.3. Specialization for configuration spaces in the ordered case
Injectivity on the left follows from the center-freeness of $\widehat{P}_{g,n}$, which can be proved using [Anderson1974] and [Birman1974, Lem. 4.2.2]. Conjugation in $\pi_{1}(\mathcal{M}_{g,n/L})$ gives a continuous action on $\widehat{P}_{g,n}$, which induces a continuous action on $P_{g,n}^{{(\ell)}}$. Consequently, there is a natural homomorphism
@@
The left vertical maps are the specialization isomorphisms defined in §3.1. The middle and right vertical maps are isomorphisms by the proof of [wat_rational_points_inp, Thm. 5.6].
It remains to justify exactness of the top row. By Lemma 4.3, the conjugation action of $\pi_{1}(\mathcal{M}_{g,n/\bar{K}})$ on $\pi_{1}(F(X_{\bar{\eta}},n))$ induces an action of $\pi_{1}(\mathcal{M}_{g,n/\bar{K}})^{{\mathrm{rel}(\ell)}}$ on $\pi_{1}(F(X_{\bar{\eta}},n))^{{(\ell)}}$. Let $W_{\bullet}$ denote the weight filtration on $\pi_{1}(F(X_{\bar{\eta}},n))^{{(\ell)}}$ constructed in [NTU]. Since the associated graded Lie algebra $\operatorname{Gr}^{W}_{\bullet}\pi_{1}(F(X_{\bar{\eta}},n))^{{(\ell)}}$ has trivial center and the filtration $W_{\bullet}$ is separated ([NTU, Cor. 2.6, p. 201]), it follows that $\pi_{1}(F(X_{\bar{\eta}},n))^{{(\ell)}}$ is center-free. Consequently, the induced homomorphism
@@ 4.4. Homotopy exact sequences in the unordered case
For $n\geq 2$ and $\ell$ an odd prime, the pro-$\ell$ completion of $S_{n}$ is trivial. Since relative pro-$\ell$ completion is right exact [HainMatsumoto2009, Prop. 2.4], applying it to (14) yields an exact sequence
@@ 5. Continuous Relative Completion
The continuous relative completion is the key tool used in the prequel to this paper [LW25]. It is the profinite analogue of the relative completion of a discrete group. Detailed introductions to the definition and properties of relative completion can be found in [hain_completion, hain_hodge_rel, hain_relwtfilt].
@@
These Lie algebras carry natural weight filtrations, inherited from the corresponding relative completions over $\mathbb{Q}$} constructed using Hodge-theoretic methods. We use only the induced structural properties of these filtrations (cf. [hain_hodge_rel, Thm. 13.1]).
@@ 5.4. Continuous relative completion in the unordered case
By the cohomological description of relative completion [hain_relwtfilt, Thm. 3.8], there is a natural isomorphism
@@
Again using [hain_relwtfilt, Thm. 3.8], we have natural isomorphisms
@@ 5.5.1. Key exact sequences for the universal curve
For the bottom row, the unordered case, right exactness of the functor $H_{1}$ implies exactness at the right; it remains only to show that the map $H_{1}(\mathfrak{p}_{g})\to H_{1}(\mathfrak{u}_{\mathcal{C}_{g,[n]},\bar{\xi}})$ is injective. To this end, we consider the natural weight filtration $W_{\bullet}\mathfrak{p}_{g}$ arising from the weighted completion of $\pi_{1}(\mathcal{C}_{g,n/\mathbb{F}_{p}})$ (see [wat_rational_points_inp]), a variant of relative completion developed by Hain and Matsumoto (see [hain-matsumoto, hain_rational]).
@@ 5.5.2. Key exact sequences for the family $\mathcal{M}_{g,n}\to\mathcal{M}_{g}$
Next, since ${\mathfrak{p}}^{\mathrm{an}}_{g,n}$ has trivial center [NTU, p. 201] and ${\mathfrak{p}}^{\mathrm{an}}_{g,n}\cong{\mathfrak{p}}_{g,n}$, the same holds for ${\mathfrak{p}}_{g,n}$ and for $F^{\mathrm{un}}_{g,n}$.
@@
Here, the vertical isomorphisms are induced by the comparison isomorphisms provided by Proposition 4.4 and Proposition 5.2. A sketch of the argument for the exactness of the bottom row at the left is as follows. The Lie algebras $\mathfrak{p}^{\mathrm{an}}_{g,n}$, $\mathfrak{u}_{g,n}$, and $\mathfrak{u}_{g}$ carry natural weight filtrations $W$ induced by their underlying Mixed Hodge Structures [hain_infpretor, Sec. 3]. Since the first homology $H_{1}$ is pure of weight 1-1, these weight filtrations coincide with the lower central series [hain_infpretor, Lem. 4.7]. The map $\mathfrak{p}^{\mathrm{an}}_{g,n}\to\mathfrak{u}_{g,n}$ is injective because $\mathfrak{p}^{\mathrm{an}}_{g,n}$ is center-free. Since the functor $\mathrm{Gr}^{W}_{\bullet}$ is exact on the category of MHS, this injectivity is preserved on the graded pieces. In particular, the induced map on the weight 1-1 component,
@@
Suppose for the sake of contradiction that $j$ is the zero map. Then $\mathrm{im}\,j=0$, which implies $\ker\pi=0$. Consequently, $\pi$ would be an isomorphism. However, this implies that $\mathrm{Hom}_{\mathrm{Sp}}^{\mathrm{cts}}(H_{1}(\mathfrak{u}_{g,[n],\bar{\xi}}),H_{\mathbb{Q}_{\ell}})=0$ by Proposition 5.2 and [hain_rational, Thm. 9.11]. This contradicts Lemma 5.3. Thus, $j$ must be an injection. This completes the proof. ∎

### v3

@@ 1. Introduction
Here $\pi_{1}(C,\bar{y})^{{(\ell)}}$ denotes the pro-$\ell$ completion of $\pi_{1}(C,\bar{y})$, and $\pi_{1}^{\prime}\!\bigl({\mathcal{C}}_{g,[n]/\bar{k}},\bar{y}\bigr)$ is a canonical quotient of $\pi_{1}\!\bigl({\mathcal{C}}_{g,[n]/\bar{k}},\bar{y}\bigr)$ obtained by the SGA 1 construction (cf. [GR71, Exp. XIII, §4]). We briefly review this construction in §2.1.
@@
This paper is a sequel to [LW25], where the corresponding nonsplitting results for universal curves with unordered marked points were established in characteristic 0. Our aim here is to extend those results to positive characteristic. Although the overall strategy follows the same general framework as in the characteristic 0 case, carrying out this extension requires substantial additional work. In characteristic 0, the essential geometric information needed for the argument is encoded in the weight filtration on relative and continuous relative completions. In positive characteristic, while continuous relative completion remains available for algebraic fundamental groups, transferring the corresponding structural properties across characteristics is far from automatic. To address this, we make systematic use of relative pro-$\ell$ completion and continuous relative completion, together with comparison isomorphisms between characteristic 0 and characteristic $p$ established in [WAT19], which allow us to transport the necessary monodromy and homotopy-theoretic information into characteristic $p$. These tools play a central role throughout the paper and are indispensable for the arguments that follow. Familiarity with the methods developed in [LW25] will therefore be beneficial to the reader.
@@ 2. Moduli of Curves with unordered marked points
The symmetric group $S_{n}$ defines a constant finite étale group scheme over $\operatorname{Spec}{\mathbb{Z}}$. By base change,
@@
is a constant finite étale group scheme over $k$ acting on ${\mathcal{M}}_{g,n/k}$ by permuting the marked points. Since ${\mathcal{M}}_{g,n/k}$ is a smooth Deligne–Mumford stack, the quotient stack
@@
is also a smooth Deligne–Mumford stack over $k$. It parametrizes smooth, proper genus $g$ curves with $n$ unordered marked points.
@@ 2.1. Homotopy exact sequences in positive characteristic
In this subsection, we briefly recall the SGA 1 construction of the homotopy exact sequence for a proper smooth morphism with geometrically connected fibers, and then apply it to the ordered and unordered universal curves (cf. [GR71, SGA 1, Exp. XIII, §4]; see also [WAT19, Prop. 4.2]).
Let $L$ be a field, and let $T$ be a connected, locally noetherian Deligne–Mumford stack over $L$. Let $f:C\longrightarrow T$ be a proper smooth morphism whose geometric fibers are connected curves of genus $g\geq 2$. Fix geometric points $\bar{\zeta}$ in $T$ and $\bar{a}$ in the fiber $C_{\bar{\zeta}}$ of $f$ over $\bar{\zeta}$. Let $\ell$ be a prime number different from $\operatorname{char}(k)$. Following [GR71, SGA 1, Exp. XIII, §4.3–4.4], set
@@
We also write $\pi_{1}(C_{\bar{\zeta}},\bar{a})^{(\ell)}$ for the maximal pro-$\ell$ quotient of $\pi_{1}(C_{\bar{\zeta}},\bar{a})$. With this notation, [WAT19, Prop. 4.2] yields the exact sequence
@@ 3.1. The pro-$\ell$ completion of $\pi_{1}(F(C,n))$
Let $K$ be the fraction field of $R$ and $\bar{K}$ an algebraic closure of $K$. For an integer $n\geq 1$, let $F(X,n)$ denote the ordered configuration space of $X$ with $n$ labeled points. The Fulton–MacPherson compactification [FM94] of $F(X,n)$ is proper and smooth over $R$, and $F(X,n)$ is the complement of a relative normal crossings divisor.
@@
for each prime $\ell\neq p$. By [GR71, Exposé XIII, Thm. 2.10, Cor. 2.9] and the fact that the Fulton–MacPherson compactification is smooth over $R$ and that $F(X,n)$ is the complement of a relative normal crossings divisor, this specialization map $\mathrm{sp}_{\ell}$ is an isomorphism.
@@ 3.2. The $\ell$-adic unipotent completion
We now pass to the $\ell$-adic unipotent completions. Following [HM03, Appendix A], let $F_{g,n}^{\mathrm{un}}$ and $F_{g,[n]}^{\mathrm{un}}$ denote the $\ell$-adic unipotent completions of $F_{g,n}$ and $F_{g,[n]}$ over $\mathbb{Q}$$_{\ell}$, respectively. These are pro-unipotent algebraic $\mathbb{Q}$$_{\ell}$}-groups equipped with Zariski-dense continuous homomorphisms
@@ 4. Relative Pro-$\ell$ Completion
To overcome this, one instead considers the relative pro-$\ell$ completion. This construction retains essential information about the mapping class group by taking into account the natural $\ell$-adic monodromy representation $\rho:{\Gamma}_{g}\longrightarrow\operatorname{Sp}_{2g}({\mathbb{Z}}_{\ell})$ and completing ${\Gamma}_{g}$ in a manner that is compatible with $\rho$. In this way, the resulting relative pro-$\ell$ completion preserves the $\ell$-adic structure coming from the geometry of ${\mathcal{M}}_{g}$. As shown in [HM09], the resulting relative pro-$\ell$ completion is nontrivial and captures the geometric structure of ${\mathcal{M}}_{g}$ much more effectively than the ordinary pro-$\ell$ completion. For instance, when marked points are present, the mapping class group embeds into its relative pro-$\ell$ completion ([HM09, Prop. 3.1]).
Because the second author applied this approach in the positive characteristic setting in [WAT19], we briefly recall below the definition and the basic properties of relative pro-$\ell$ completion. Full details and extensive background can be found in [HM09].
@@ 4.1. The relative pro-$\ell$ completion of the mapping class group
When $\operatorname{char}(L)=p>0$ and $\ell\neq p$, the second author proved in [WAT19, Thm. 5.6] that the relative pro-$\ell$ completion of $\pi_{1}({\mathcal{M}}_{g,n/\bar{L}},\bar{\zeta})$ with respect to $\rho_{{\mathbb{Z}}_{\ell},\bar{\zeta}}$ is naturally isomorphic to ${\Gamma}_{g,n}^{{\mathrm{rel}(\ell)}}$, the relative pro-$\ell$ completion of the mapping class group ${\Gamma}_{g,n}$ with respect to its standard symplectic representation
@@ 4.2. Specialization for the universal curve in the ordered case
The rows are exact by right exactness of relative pro-$\ell$ completion and the center-freeness of $\pi_{1}(C_{\bar{\xi}})^{{(\ell)}}(\cong\pi_{1}(C_{\bar{\eta}})^{{(\ell)}})$, and the right vertical homomorphisms are isomorphisms (see the proof of [WAT19, Thm. 5.6]). It therefore follows that all middle vertical homomorphisms are isomorphisms. ∎
@@ 4.3. Specialization for configuration spaces in the ordered case
Injectivity on the left follows from the center-freeness of $\widehat{P}_{g,n}$, which can be proved using [AND74] and [BIR74, Lem. 4.2.2]. Conjugation in $\pi_{1}(\mathcal{M}_{g,n/L})$ gives a continuous action on $\widehat{P}_{g,n}$, which induces a continuous action on $P_{g,n}^{{(\ell)}}$. Consequently, there is a natural homomorphism
@@
The left vertical maps are the specialization isomorphisms defined in §3.1. The middle and right vertical maps are isomorphisms by the proof of [WAT19, Thm. 5.6].
It remains to justify exactness of the top row. By Lemma 4.3, the conjugation action of $\pi_{1}(\mathcal{M}_{g,n/\bar{K}})$ on $\pi_{1}(F(X_{\bar{\eta}},n))$ induces an action of $\pi_{1}(\mathcal{M}_{g,n/\bar{K}})^{{\mathrm{rel}(\ell)}}$ on $\pi_{1}(F(X_{\bar{\eta}},n))^{{(\ell)}}$. Let $W_{\bullet}$ denote the weight filtration on $\pi_{1}(F(X_{\bar{\eta}},n))^{{(\ell)}}$ constructed in [NTU95]. Since the associated graded Lie algebra $\operatorname{Gr}^{W}_{\bullet}\pi_{1}(F(X_{\bar{\eta}},n))^{{(\ell)}}$ has trivial center and the filtration $W_{\bullet}$ is separated ([NTU95, Cor. 2.6, p. 201]), it follows that $\pi_{1}(F(X_{\bar{\eta}},n))^{{(\ell)}}$ is center-free. Consequently, the induced homomorphism
@@ 4.4. Homotopy exact sequences in the unordered case
For $n\geq 2$ and $\ell$ an odd prime, the pro-$\ell$ completion of $S_{n}$ is trivial. Since relative pro-$\ell$ completion is right exact [HM09, Prop. 2.4], applying it to (14) yields an exact sequence
@@ 5. Continuous Relative Completion
The continuous relative completion is the key tool used in the prequel to this paper [LW25]. It is the profinite analogue of the relative completion of a discrete group. Detailed introductions to the definition and properties of relative completion can be found in [HAI93, HAI98, HAI08].
@@
These Lie algebras carry natural weight filtrations, inherited from the corresponding relative completions over $\mathbb{Q}$} constructed using Hodge-theoretic methods. We use only the induced structural properties of these filtrations (cf. [HAI98, Thm. 13.1]).
@@ 5.4. Continuous relative completion in the unordered case
By the cohomological description of relative completion [HAI08, Thm. 3.8], there is a natural isomorphism
@@
Again using [HAI08, Thm. 3.8], we have natural isomorphisms
@@ 5.5.1. Key exact sequences for the universal curve
For the bottom row, the unordered case, right exactness of the functor $H_{1}$ implies exactness at the right; it remains only to show that the map $H_{1}(\mathfrak{p}_{g})\to H_{1}(\mathfrak{u}_{\mathcal{C}_{g,[n]},\bar{\xi}})$ is injective. To this end, we consider the natural weight filtration $W_{\bullet}\mathfrak{p}_{g}$ arising from the weighted completion of $\pi_{1}(\mathcal{C}_{g,n/\mathbb{F}_{p}})$ (see [WAT19]), a variant of relative completion developed by Hain and Matsumoto (see [HM03, HAI11]).
@@ 5.5.2. Key exact sequences for the family $\mathcal{M}_{g,n}\to\mathcal{M}_{g}$
Next, since ${\mathfrak{p}}^{\mathrm{an}}_{g,n}$ has trivial center [NTU95, p. 201] and ${\mathfrak{p}}^{\mathrm{an}}_{g,n}\cong{\mathfrak{p}}_{g,n}$, the same holds for ${\mathfrak{p}}_{g,n}$ and for $F^{\mathrm{un}}_{g,n}$.
@@
Here, the vertical isomorphisms are induced by the comparison isomorphisms provided by Proposition 4.4 and Proposition 5.2. A sketch of the argument for the exactness of the bottom row at the left is as follows. The Lie algebras $\mathfrak{p}^{\mathrm{an}}_{g,n}$, $\mathfrak{u}_{g,n}$, and $\mathfrak{u}_{g}$ carry natural weight filtrations $W$ induced by their underlying Mixed Hodge Structures [HAI97, Sec. 3]. Since the first homology $H_{1}$ is pure of weight 1-1, these weight filtrations coincide with the lower central series [HAI97, Lem. 4.7]. The map $\mathfrak{p}^{\mathrm{an}}_{g,n}\to\mathfrak{u}_{g,n}$ is injective because $\mathfrak{p}^{\mathrm{an}}_{g,n}$ is center-free. Since the functor $\mathrm{Gr}^{W}_{\bullet}$ is exact on the category of MHS, this injectivity is preserved on the graded pieces. In particular, the induced map on the weight 1-1 component,
@@
Suppose for the sake of contradiction that $j$ is the zero map. Then $\mathrm{im}\,j=0$, which implies $\ker\pi=0$. Consequently, $\pi$ would be an isomorphism. However, this implies that $\mathrm{Hom}_{\mathrm{Sp}}^{\mathrm{cts}}(H_{1}(\mathfrak{u}_{g,[n],\bar{\xi}}),H_{\mathbb{Q}_{\ell}})=0$ by Proposition 5.2 and [HAI11, Thm. 9.11]. This contradicts Lemma 5.3. Thus, $j$ must be an injection. This completes the proof. ∎